# SLIM: Scalable and Lightweight LiDAR Mapping in Urban Environments

### v1

@@ IV-A2 Vectorized Mapping
Constructing line and plane observations is merely the first step. The crucial step is to parameterize the observations into landmarks for long-term operations. To decrease the problem scale, we sample dense robot poses from odometry as keyframes. Regarding the landmarks, we define a line landmark as $\mathbf{l}_{\mathcal{L}}:=\langle l_{l},\mathbf{c}_{\mathcal{L}},\mathbf{n}_{\mathcal{L}},\mathbf{p}_{\mathcal{L}},\left\{\mathbf{f}_{{\mathcal{L}}_{i}}\right\}\rangle$ and a plane landmark as $\mathbf{l}_{\mathcal{S}}:=\langle l_{s},\mathbf{c}_{\mathcal{S}},\mathbf{n}_{\mathcal{S}},\mathbf{p}_{\mathcal{S}},\left\{\mathbf{f}_{{\mathcal{S}}_{i}}\right\}\rangle$, which include a label, centroid, normal, minimum parameter block, and their observations across different keyframes. Specifically, the normal is a direction vector for a line and a normal vector for a plane. The observations encodes that the landmark is associated with a specific keyframes, and this information can be obtained through centroid-based nearest neighbor searching when applying LiDAR odometry methods.
@@
We have constructed a factor graph structure similar to visual SLAM. The next step is to parameterize the landmarks to enable optimization methods for refining the results. Meanwhile, a minimum parameter block is needed to reduce the problem scale as mentioned. The point-normal form, i.e., the normal vectors $\mathbf{n}$ and random points $\mathbf{c}$, is over-parameterized for lines and planes and unsuitable for optimization problems. To address this, we propose representing an infinite line as $\mathbf{p}_{\mathcal{L}}:=\langle\alpha,\beta,x,y\rangle\in\mathbb{R}^{4}$, where $\alpha$ and $\beta$ represent the direction, and $x$ and $y$ represent the offset translation on the $xOy$ plane. Similarly, an infinite plane is formulated as $\mathbf{p}_{\mathcal{S}}:=\langle\alpha,\beta,d\rangle\in\mathbb{R}^{3}$, where $\alpha$ and $\beta$ represent the direction, and $d$ represents the offset translation on the $z$-axis, thus formulating the desired minimum parameter blocks. Figure 3 illustrates how we obtain arbitrary lines and planes using our minimal parameters.
@@
Figure 3: Formulation of line and plane landmarks. We can obtain $\mathbf{l}(x,y)$ by translating $(x,y)$, and then $\mathbf{l}(\mathbf{R}(\alpha,\beta),x,y)$ using a two-degree-of-freedom rotation $\mathbf{R}(\alpha,\beta)$. Similarly, we can obtain $\mathbf{s}(d)$ by translating $d$, and then achieve $\mathbf{s}(\mathbf{R}(\alpha,\beta),d)$ by the rotation. For any line or plane, a corresponding transformation always exists to formulate representation.
@@
Theoretically, all infinite lines and infinite planes could be transformed from the original line $\mathbf{u}_{z}$ and original plane $xOy$ in two steps, where $\mathbf{u}_{x},\mathbf{u}_{y},\mathbf{u}_{z}$ are unit vectors along $x$-axis, $y$-axis and $z$-axis. or line landmarks, we first apply an offset $(x,y)$ to the original line $\mathbf{u}_{z}$ to obtain the translated line $\mathbf{l}(x,y)$. Then, we rotate the nearest point $(x,y)$ using one 2 degrees-of-freedom (DoF) rotation matrix $\mathbf{R}(\alpha,\beta)$ to obtain the final line $\mathbf{l}(\mathbf{R}(\alpha,\beta),x,y)$. The 2 DoF rotation matrix $\mathbf{R}(\alpha,\beta)$ is defined as follows:
@@
For plane landmarks, the original plane $xOy$ is translated along the $z$ axis and rotated using a 2 DoF rotation matrix, resulting in the final plane $\mathbf{s}(\mathbf{R}(\alpha,\beta),d)$.
@@ IV-B1 Map Merge Strategy
Typical global map merge methods first transform full laser scans into handcrafted or learning-based global descriptors for place retrieval. Then, point cloud registration is indispensable for relative transformation estimation, as introduced in Section II. Existing approaches often require extensive raw information, such as the original point cloud or embedding feature maps, leading to significant data demands for feature learning and transmission. Additionally, these methods mostly focus on scan-to-scan place retrieval rather than utilizing multi-frame information to represent a place. Single scan-based methods are more sensitive to displacement and lack robustness for map merging [21].
@@
To address this issue, we design a two-step method to reduce the problem size. First, we cluster the plane landmarks that lie on the same infinite plane into a single plane, particularly for those on large buildings and road surfaces. The line landmarks are maintained without clustering. Second, we partition the base map into blocks $\left\{\mathcal{G}{b_{p}}\right\}$ and the submap into blocks $\left\{\mathcal{G}{s_{q}}\right\}$ along the robot trajectory, instead of directly matching the submap to the base map. Each block comprises a host keyframe and a set of landmarks around it. These two steps make the map structure more compact for finding the maximum clique by reducing the number of nodes in a block, thereby leading to more efficient block registration for multi-session map merging.
@@ IV-B2 Pairwise Block Global Registration
Figure 4: Two cases of the block registration. We reduce the number of plane landmarks by clustering, thus speeding up the registration process for map merge. The lines between the clustered landmarks represent the data associations (correspondences) between the block in different coordinates.
We denote all line and plane correspondences as $\mathcal{C}_{\mathcal{L}}=\{(k,l)\}$ and $\mathcal{C}_{\mathcal{S}}=\{(k,l)\}$, respectively, where $k$ and $l$ are the indices of landmarks in blocks $\mathcal{G}_{b_{p}}$ and $\mathcal{G}_{s_{q}}$ (blocks belong to the base map and the submap, respectively). The pairwise compatibility test is performed for any two correspondences as follows:
@@
where $(k_{i},l_{i})$ and $(k_{j},l_{j})\in\mathcal{C}_{\mathcal{L}}\cup\mathcal{C}_{\mathcal{S}}$. Since the Grassmannian metric is $\text{SE}(3)$-transformation invariant, we define a truncation threshold $\delta$ to determine whether both correspondences pass the compatibility test. If passed, an edge is connected between these correspondences in the compatibility graph. The maximum clique is then found using a fast exact parallel finder algorithm [82], where the nodes in this clique represent the putative line and plane correspondence inlier sets $\mathcal{C}{\mathcal{L}}^{*}$ and $\mathcal{C}{\mathcal{S}}^{*}$. Figure 4 presents two cases of block registration for map merge in this study.
With the pruned correspondence results, we can establish a hybrid registration utilizing lines and planes to estimate the relative poses between graphs (blocks), which is formulated as follows:
@@
where $\mathbf{c}=b$, $\mathbf{n}=A$ for lines and $\mathbf{n}\in\text{span}(A)^{\perp}$ for planes, respectively. The term $\mathbf{T}^{b_{i}}_{s_{i}}$ is the relative transformation between block $\mathcal{G}_{b_{i}}$ and block $\mathcal{G}_{s_{i}}$; $\rho(\cdot)$ denotes a robust kernel function, which further reduces the influence of potential correspondence outliers resulting from an inappropriate selection of the threshold $\delta$, as discussed in [80]. At the start of the iteration, the initial value of $\mathbf{T}^{b_{i}}_{s_{i}}$ is set to the identity transformation.
@@ IV-B3 Loop Outlier Rejection
where $\mathbf{\hat{T}}^{s_{l}}_{s_{k}}$ and $\mathbf{\hat{T}}^{b_{k}}_{b_{l}}$ are from odometry trajectories in different sessions. If $C(\mathbf{T}^{b_{k}}_{s_{k}},\mathbf{T}^{b_{l}}_{s_{l}})$ are small enough, the pair of relative pose $\mathbf{T}^{b_{k}}_{s_{k}}$ and $\mathbf{T}^{b_{l}}_{s_{l}}$ are considered to be consistent. Similarly, the problem of solving the internal-consistent set of relative poses is also a maximum clique problem [82]. The pruned set of relative poses (transformations) is denoted as $\mathcal{L}_{\mathcal{O}}=\left\{\mathbf{T}^{b_{i}}_{s_{i}}\right\}$, which will be used for subsequent PGO, thus significantly reducing the drifts by odometry.
@@ IV-C1 Pose Graph Optimization
where $w$ represents the world coordinate. Despite the relative poses between blocks, the LiDAR odometry could also provide rigid transformations between consecutive keyframe poses, as illustrated in Section IV-A1. The odometry-based residual function is defined as
@@
where $\mathcal{O}_{b}$ and $\mathcal{O}_{s}$ are the set of all odometry residuals of the base map and submap. $\mathbf{\Lambda}_{\mathcal{O}}$ and $\mathbf{\Lambda}_{\mathcal{L}_{\mathcal{O}}}$ are tunable parameters to describe the uncertainties. Empirically, $\mathbf{\Lambda}_{\mathcal{O}}$ is determined by the drift rate of applied odometry; $\mathbf{\Lambda}_{\mathcal{L}_{\mathcal{O}}}$ is insensitive to different datasets or scenarios, hence, a general threshold is acceptable. After solving the PGO, odometry drift will be preliminarily eliminated, and all keyframes of multi-session data will be aligned more closely in the world coordinate. More specifically, We will merge the base map $\mathcal{M}_{b}$ and submap $\mathcal{M}_{s}$ at the data structure level, including the set of keyframes, the set of odometry residuals, landmarks, and all observations.
PGO enhances the accuracy of poses, thereby improving overall mapping precision through pose adjustments. However, in some cases, the optimized results become unsatisfactory due to significant odometry drift. The primary issue is that drifted odometry leads to incorrect map vectorization, resulting in inaccurate relative pose calculations between sessions. To address this, we propose a BA approach to enhance map refinement by jointly adjusting landmarks and keyframe poses.
@@ IV-C2 Bundle Adjustment
Inspired by visual BA [84], we design the residual function between the observation in the keyframe and the corresponding landmark in the world. Geometrically, a line can be defined by two points. Thus, every line landmark has two point-to-infinite-line residuals:
@@
To this end, all multi-session robot poses and landmarks are merged and optimized jointly, as detailed in Sections IV-B and IV-C. It is worth noting that these two modules utilize only parameterized lines and planes provided by the map vectorization in Section IV-A. Using such lightweight representations ensures memory efficiency while maintaining mapping accuracy. The experimental sections will validate the superiority of the proposed modules individually in terms of accuracy, efficiency, and generalizability across different datasets.
V Map-centric Nonlinear Factor Recovery
The dimensions of optimization will increase over time, which is a typical issue for long-term mapping. This growth leads to significant time costs for optimization. This section presents a map-centric NFR that aims to maintain a factor graph with sparsified poses and unaltered landmarks as sessions increase. Specifically, the objective is to marginalize redundant keyframe poses while preserving the original problem structure for BA. The proposed map-centric NFR maintains global consistency while bounding the computational requirements for long-term mapping
@@ V-A Motivation and Problem Formulation
In classical SLAM systems, marginalization is a crucial technique designed to prune the graph and preserve prior information without breaking the observability of the system [85]. The theoretical basis for this technique is the Schur complement [86]. However, the Schur complement generally introduces a ”fill-in” of non-zero blocks in the Hessian matrix, significantly slowing down the nonlinear optimization. Nonlinear factor recovery [73, 74] approximates a new distribution of the original problem with fewer recovered factors. The mathematical description of the NFR problem will be detailed as follows.
@@
Figure 6 presents a graphical illustration of NFR and its corresponding graph topology. Specifically, map-centric NFR encompasses two primary steps: first, a new topology is reconstructed from the original by keyframe removal; second, the information matrices of these residuals, which also serve as the solutions to the KLD problem, will be calculated as described in Equation (18). These steps will be detailed in the following section.
@@ V-B1 Topology Reconstruction
    All retained keyframes are connected by a minimum spanning tree, and $N-1$ keyframe-to-keyframe residuals are preserved, where $N$ is the number of retained keyframes.
@@
The reconstructed Jacobian matrix is always square and invertible, meeting the requirements for Equation (18). This is due to two reasons: the dimension of all keyframe-to-landmark residuals always equals the dimension of landmark states; the sum of the dimensions of keyframe-to-keyframe residuals and the dimension of the prior keyframe pose residual exactly equals the dimension of retained keyframe states.
Keyframe removal leads to information loss; hence, marginalization is needed to maintain the necessary information for mapping. The classical Schur-complement-based marginalization is formulated as follows. Assume that we have $M$ line landmarks, $N$ plane landmarks, $P$ retained keyframes and $Q$ marginalized keyframes. The dimensions are: retained state variables $r=M\times 4+N\times 3+P\times 6$; landmarks $l=M\times 4+N\times 3$; retained frame $n=P\times 6$; marginalized state variables $m=Q\times 6$. Schur-complement-based marginalization is conducted when redundant keyframes are selected, providing the marginalized prior Hessian matrix. Given the Hessian matrix before marginalization $\mathbf{H}$, the covariance matrix $\mathbf{\Sigma}_{rr}^{m}$ after marginalization is derived as:

### v2

@@ IV-A2 Vectorized Mapping
Parameterizing the observations into landmarks is crucial for long-term maintenance. To reduce the complexity, we sample dense robot poses from odometry as keyframes. Regarding the landmarks, we define a line landmark as $\mathbf{l}_{\mathcal{L}}:=\langle l_{l},\mathbf{c}_{\mathcal{L}},\mathbf{n}_{\mathcal{L}},\mathbf{p}_{\mathcal{L}},\left\{\mathbf{f}_{{\mathcal{L}}_{i}}\right\}\rangle$ and a plane landmark as $\mathbf{l}_{\mathcal{S}}:=\langle l_{s},\mathbf{c}_{\mathcal{S}},\mathbf{n}_{\mathcal{S}},\mathbf{p}_{\mathcal{S}},\left\{\mathbf{f}_{{\mathcal{S}}_{i}}\right\}\rangle$, which include a label, centroid, normal, minimum parameter block, and their observations across different keyframes. Specifically, the normal is a directional vector for a line and a normal vector for a plane. The observations encodes that the landmark is associated with a specific keyframes, and this association can be obtained through centroid-based nearest neighbor searching when applying LiDAR odometry. Essentially, we have constructed a factor graph structure similar to visual SLAM.
@@
The next step is to parameterize the landmarks to enable optimization modules for map refinement. Meanwhile, a minimum parameter block is needed to reduce the problem scale as mentioned. The point-normal form, i.e., the normal vectors $\mathbf{n}$ and random points $\mathbf{c}$, is over-parameterized for lines and planes and unsuitable for optimization problems. To address this, we propose representing an infinite line as $\mathbf{p}_{\mathcal{L}}:=\langle\alpha,\beta,x,y\rangle\in\mathbb{R}^{4}$, where $\alpha$ and $\beta$ represent the direction, and $x$ and $y$ represent the offset translation on the $xOy$ plane. Similarly, an infinite plane is formulated as $\mathbf{p}_{\mathcal{S}}:=\langle\alpha,\beta,d\rangle\in\mathbb{R}^{3}$, where $\alpha$ and $\beta$ represent the direction, and $d$ represents the offset translation on the $z$-axis, thus formulating the desired minimum parameter blocks.
@@
Figure 3: Formulation of line and plane landmarks. On the left side, the line $\mathbf{l}(x,y)$ can be obtained through a translation $(x,y)$, followed by the derivation of $\mathbf{l}(\mathbf{R}(\alpha,\beta),x,y)$ using a 2-DoF rotation $\mathbf{R}(\alpha,\beta)$. Similarly, on the right side, the plane $\mathbf{s}(d)$ is formulated by applying a translation $d$, and subsequently transformed into $\mathbf{s}(\mathbf{R}(\alpha,\beta),d)$ via rotation. A corresponding transformation always exists for any line or plane to formulate landmarks.
@@
Figure 3 illustrates how we obtain arbitrary lines and planes using our minimal parameters. Theoretically, all infinite lines and infinite planes could be transformed from the original line $\mathbf{u}_{z}$ and the original plane $xOy$ in two steps, where $\mathbf{u}_{x},\mathbf{u}_{y},\mathbf{u}_{z}$ are unit vectors along $x$-axis, $y$-axis and $z$-axis. For line landmarks, we first apply an offset $(x,y)$ to the original line $\mathbf{u}_{z}$ to obtain the translated line $\mathbf{l}(x,y)$. Then, we rotate the nearest point $(x,y)$ using one 2 DoF rotation matrix $\mathbf{R}(\alpha,\beta)$ to obtain the final line $\mathbf{l}(\alpha,\beta,x,y)$. The 2 DoF rotation matrix $\mathbf{R}(\alpha,\beta)$ is defined as follows:
@@
For plane landmarks, the original plane $xOy$ is translated along the $z$ axis and rotated using a 2 DoF rotation matrix, resulting in the final plane $\mathbf{s}(\alpha,\beta,d)$.
@@
Detailed residual formulations, utilizing the proposed line and plane representations, are presented in Appendix VIII. The optimization process (map refinement) is introduced in Section IV-C and further illustrated in Appendix VIII.
@@ IV-B1 Map Merge Strategy
Typical global map merge methods first transform full laser scans into handcrafted or learning-based global descriptors for place retrieval. Then, point cloud registration is indispensable for relative transformation estimation, as introduced in Section II. Existing approaches often require extensive raw information for place recognition, such as the original point cloud or embedding feature maps, leading to significant data demands for learning and transmission. Additionally, these methods mostly focus on scan-to-scan place retrieval rather than utilizing multi-frame information to encode a place. Single scan-based methods are more sensitive to displacement and lack robustness [20].
@@
To address this issue, we design a two-step method to reduce the problem size. First, we cluster the plane landmarks that lie on the same infinite plane into a single plane, particularly for those on large buildings and road surfaces. The line landmarks are maintained without clustering. Second, we partition the base map into blocks $\left\{\mathcal{G}{b_{p}}\right\}$ and the submap into blocks $\left\{\mathcal{G}{s_{q}}\right\}$ along the robot trajectory, instead of directly matching the submap to the base map. Each block comprises a host keyframe and a set of landmarks around it. These two steps make the map structure more compact for finding the maximum clique by reducing the number of nodes in a block, thereby leading to more efficient block registration.
@@ IV-B2 Pairwise Block Global Registration
Figure 4: Two registration cases at different places, using the proposed block registration method. We reduce the number of plane landmarks by clustering, thus speeding up the registration process for map merge. The lines between the clustered landmarks represent the data associations (correspondences) between the block in different coordinates.
We denote all line and plane correspondences as $\mathcal{C}_{\mathcal{L}}=\{(k,l)\}$ and $\mathcal{C}_{\mathcal{S}}=\{(k,l)\}$, where $k$ and $l$ are the indices of landmarks in blocks $\mathcal{G}_{b_{p}}$ and $\mathcal{G}_{s_{q}}$ (blocks belong to the base map and the submap, respectively). The pairwise compatibility test for any two correspondences is formulated as follows:
@@
where $(k_{i},l_{i})$ and $(k_{j},l_{j})\in\mathcal{C}_{\mathcal{L}}\cup\mathcal{C}_{\mathcal{S}}$. Since the Grassmannian metric is $\text{SE}(3)$-transformation invariant, we define a truncation threshold $\delta$ to determine whether both correspondences pass the compatibility test. If passed, an edge is connected between these correspondences in the compatibility graph. The maximum clique is then found using a fast exact parallel finder algorithm [80], where the nodes in this clique represent the putative line and plane correspondence inlier sets $\mathcal{C}_{\mathcal{L}}^{*}$ and $\mathcal{C}_{\mathcal{S}}^{*}$. Figure 4 presents two cases of block registration for map merge in this study.
With the pruned correspondence results, we can establish a hybrid registration to estimate the relative poses between graphs (blocks), which is formulated as follows:
@@
where $\mathbf{c}=b$, $\mathbf{n}=A$ for lines and $\mathbf{n}\in\text{span}(A)^{\perp}$ for planes. The term $\mathbf{T}^{b_{i}}_{s_{i}}$ is the relative transformation between block $\mathcal{G}_{b_{i}}$ and block $\mathcal{G}_{s_{i}}$; $\rho(\cdot)$ denotes a robust kernel function, which further reduces the influence of potential correspondence outliers resulting from an inappropriate selection of the threshold $\delta$, as discussed in [78]. At the start of iterations, the initial value of $\mathbf{T}^{b_{i}}_{s_{i}}$ is set to the identity transformation.
@@ IV-B3 Loop Outlier Rejection
where $\mathbf{\hat{T}}^{s_{l}}_{s_{k}}$ and $\mathbf{\hat{T}}^{b_{k}}_{b_{l}}$ are from odometry trajectories in different sessions. If $C(\mathbf{T}^{b_{k}}_{s_{k}},\mathbf{T}^{b_{l}}_{s_{l}})$ are small enough, the pair of relative pose $\mathbf{T}^{b_{k}}_{s_{k}}$ and $\mathbf{T}^{b_{l}}_{s_{l}}$ are considered to be consistent. Similarly, the problem of solving the internal-consistent set of relative poses is also a maximum clique problem [80]. The pruned set of relative poses (transformations) is denoted as $\mathcal{L}_{\mathcal{O}}=\left\{\mathbf{T}^{b_{i}}_{s_{i}}\right\}$, and will be used for subsequent PGO.
@@ IV-C1 Pose Graph Optimization
where $w$ represents the world coordinate. In addition to the relative poses between blocks, the LiDAR odometry could also provide rigid transformations between consecutive keyframe poses, as illustrated in Section IV-A1. The odometry-based residual function is defined as
@@
where $\mathcal{O}_{b}$ and $\mathcal{O}_{s}$ are the set of all odometry residuals of the base map and submap. $\mathbf{\Lambda}_{\mathcal{O}}$ and $\mathbf{\Lambda}_{\mathcal{L}_{\mathcal{O}}}$ are tunable parameters to describe the uncertainties. Empirically, $\mathbf{\Lambda}_{\mathcal{O}}$ is determined by the drift rate of applied odometry; $\mathbf{\Lambda}_{\mathcal{L}_{\mathcal{O}}}$ is insensitive to different datasets or scenarios, hence, a general threshold is acceptable. After solving the PGO, odometry drift will be preliminarily eliminated, and all keyframes of multi-session data will be aligned more closely in the world coordinate. More specifically, We will merge the base map $\mathcal{M}_{b}$ and submap $\mathcal{M}_{s}$ at the data structure level, including the set of keyframes, the set of odometry residuals, landmarks, and all observations. The process of merging landmarks will be detailed as follows.
[…]
@@ IV-C2 Bundle Adjustment
Drifted odometry could result in incorrect map vectorization, leading to inaccurate relative pose estimations between sessions that cannot be fully resolved by PGO. To address this, we propose a BA approach to improve map refinement by jointly adjusting landmarks and poses. Inspired by visual BA [82], we design the residual function between the observation in the keyframe and the corresponding landmark in the world. Geometrically, a line can be defined by two points. Thus, every line landmark has two point-to-infinite-line residuals:
@@
The overall map refinement is performed in two steps: a pose-only approach (PGO) followed by a pose-landmark joint approach (BA), which offers two key advantages. First, the refinement process is conducted in a coarse-to-fine manner, which enhances the convergence of optimization. Performing BA directly without PGO may lead to convergence failure, as observed during tests using the SLIM system. Second, the optimized poses obtained from PGO serve as an initial guess for BA, thus improving the efficiency of the map refinement.
To this end, multi-session maps are merged and refined sequentially, as detailed in Sections IV-B and IV-C. It is worth noting that these two modules utilize only parameterized lines and planes provided by the map vectorization in Section IV-A. Using such lightweight representations ensures memory efficiency while maintaining mapping accuracy. The experimental sections will validate the superiority of the proposed modules individually in terms of accuracy, efficiency, and generalizability across different datasets.
V Map-centric Marginalization
The dimensions of optimization will increase over time, which is a typical issue for long-term mapping. This growth leads to significant time costs for optimization. This section presents a map-centric marginalization that aims to maintain a factor graph with sparsified poses and unaltered landmarks as sessions increase. The proposed map-centric marginalization maintains global consistency while bounding the computational requirements for long-term mapping.
@@ V-A Motivation and Problem Formulation
In classical SLAM systems, marginalization is a crucial technique designed to prune the graph and preserve prior information without breaking the observability of the system [83]. The theoretical basis for this technique is the Schur complement [84]. However, the Schur complement generally introduces a “fill-in” of non-zero blocks in the Hessian matrix, significantly slowing down the nonlinear optimization. Nonlinear factor recovery [71, 72] approximates a new distribution of the original problem with fewer recovered factors. NFR is a classical method and plays a key role in the marginalization process of this study. The mathematical description of the NFR problem will be detailed as follows.
@@
Figure 6 presents a graphical illustration of NFR and its corresponding graph topology. Specifically, our map-centric marginalization encompasses two primary steps, as detailed in the following section. First, a new topology is reconstructed from the original by keyframe removal; second, the information matrices of these residuals, which also serve as the solutions to the KLD problem, will be calculated as described in Equation (20).
@@ V-B1 Topology Reconstruction
    All retained keyframes are connected by a minimum spanning tree, and $N-1$ keyframe-to-keyframe residuals are preserved ($N$ is the number of retained keyframes).
@@
The reconstructed Jacobian matrix is always square and invertible, meeting the requirements for Equation (20). This is due to two reasons: the dimension of all keyframe-to-landmark residuals always equals the dimension of landmark states; the sum of the dimensions of keyframe-to-keyframe residuals and the dimension of the prior keyframe pose residual exactly equals the dimension of retained keyframe states. It is worth noting that the topology reconstruction is performed after map refinement in Section IV-C. Therefore, even though the minimum spanning tree no longer includes any loop closures, the topology reconstruction does not affect the utilization of loop closures for map refinement.
Keyframe removal leads to information loss. Therefore, marginalization is needed to maintain the necessary information for mapping. The classical Schur-complement-based marginalization is formulated as follows. Assume that we have $M$ line landmarks, $N$ plane landmarks, $P$ retained keyframes and $Q$ marginalized keyframes. The dimensions are: retained state variables $r=M\times 4+N\times 3+P\times 6$; landmarks $l=M\times 4+N\times 3$; retained frame $n=P\times 6$; marginalized state variables $m=Q\times 6$. Schur-complement-based marginalization is conducted when redundant keyframes are selected, providing the marginalized prior Hessian matrix. Given the Hessian matrix before marginalization $\mathbf{H}$, the covariance matrix $\mathbf{\Sigma}_{rr}^{m}$ after marginalization is derived as: